# KEN KANGDA WREN1

### v7

@@ 1.3 The Economic Decomposition of Risk-Pricing under Model-Risk
where $\check{\mathbf{E}}_{{B}}[\cdot|\cdot]$ denotes expectations under its $B$-sure RNE measure $\check{\mathbf{W}}^{T}_{B}\sim{\mathbf{W}}^{T}_{B}$, leading to RNE outlook $\check{y}_{B}(T-t,t)=\int_{t}^{T}\check{r}_{B,s}ds$ as in (4), and $B$-sure risk-premium $RP^{B}(t)$ satisfying: $\forall t\leq T$,
@@
where (14) ensures consistency with asset-property Item-2, Section 1.1, while (15), the internal consistency of the pricing scheme under model-risk $B$.
@@
$k^{A}_{t}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}\geq 0,$ (19)
@@
In ambiguity-aversion settings, the primary utility function, enforcing risk-aversion, and the secondary utility/variational-control, enforcing uncertainty-aversion, can be dialled independently and arbitrarily close to neutrality (linearity), to achieve the above separation. In classical and some parameter-learning settings, with a single utility, the above obtains by replacing payoffs with their model-sure expectations perturbed by suitably small noise.
@@
Asset-pricing under model-risk $B$ is said to be usual if its price-of-$B$-risk $k^{A}_{t}$, (19), is a time-homogenous twice-differentiable function of its model-risk belief [^3] Economic risk-pricing, as the solution to some optimisation scheme, is measurable to the natural filtration of its input $(\{Z_{t}\},\{\pi^{B}_{t}(\mathbf{D}_{t})\})$; so it is at any time $t$ a function of $(Z_{t},\pi^{B}_{t}(\mathbf{D}_{t}))$ (Proposition 4.9 of Breiman (1992)). Any explicit dependence on $t$ or $Z_{t}$, if at all, is presumed here weak or non-existent, at least for $t\ll T$. $\pi^{B}_{t}$, $\forall t\ll T$.
@@ 2.2.1 The 1-Parameter Price-of-Model-Risk Formula
The difference process $\{{\pi}^{+}_{t}\}-\{{\Pi}^{+}_{t}\}=\{k^{\Pi}_{t}\sigma^{\pi}_{t}\}$ has a fixed sign. Unless the model-risk is consistent (see Item-2, Section 1.1, and (14)), viable model-risk premia $\{B\_RP_{t}\}$ (see (18)) must be negative sometimes and so correspond to no known economic risk-pricing theories.
@@ 2.2.2 Connection with Familiar Price-Dynamics
For better focus, let any $B$-sure risk-premia be $B$-independent henceforth, to be written as $\check{rp}(t)=\int_{t}^{T}\check{R}_{s}ds>0$, with $\check{R}_{t}$ denoting the $B$-sure risk-premium drift at any $t<T$; note then $Y^{\Delta}(t)=S^{\Delta}(t)=\int_{t}^{T}r^{Z\Delta}_{s}ds$ as a result. Canonical pricing (22) becomes:
@@
$dS_{t}(B)=dY^{-}_{t}(B)+S^{\Delta}(t)d\Pi^{+}_{t}(B)-r^{Z\Delta}_{t}\Pi^{+}_{j}(B)dt+\check{R}_{t}dt.$ (25)
By setup (Section 1.1) and inference dynamics (A.2-A.3), process $\{d\Pi^{+}_{t}\}$ has two sources of noise, one from $\{Z_{t}\}$, given by $\{\sigma^{lZ}_{t}dw^{Z}_{t}\}$, with $\sigma^{lZ}_{t}:=r^{Z\Delta}_{t}/\sigma^{Z}_{t}\ll 1$ (’signal-to-noise’), one from independent $B$-informative data $\{D_{t}\}$, modelled in parallel ((2)), which can be absorbed through $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$ and $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$; then:
@@
$\mu^{\Pi,\pi}_{t}:=\Sigma_{B}{\pi}^{B}_{t}{\mu^{\Pi,B}_{t}}=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot(\pi^{+}_{t}-\Pi^{+}_{t})=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot k^{\Pi}_{t}\sigma^{\pi}_{t}\geq 0;$ (29)
@@
$+$ $\big{[}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{j}(B)\big{)}\cdot r^{Z\Delta}_{t}dt\big{]}$ (31)
@@
Practical cases often see $B$-sure dynamics (30) and model-drift (31) dominated by model-inference (32): $\check{R}_{t}\ll S^{\Delta}(t)$ and $r^{Z\Delta}_{t}\ll S^{\Delta}(t)$; indeed firmed-up values $\{Z_{t}\}$ are often nil or fixed so that asset-pricing is just a matter of model-risk inference, with $\{dS_{t}\}$ reduced to the inferential test equation of Peskir and Shirayev (2006) (for constant $\{r^{D\Delta}_{t}\}$ and $\{\sigma^{D}_{t}\}$). Further, it often holds that relative variance $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}\ll 1$ and so $(\sigma^{l}_{t})^{2}\approx(\sigma^{lD}_{t})^{2}$: model-risk is the prime price-mover, not tangible value. Dispersion among such assets is thus driven mainly by their differing underlying model-risk inference processes.
@@
Expansion in (K1)(K-1), given Remark 7, recovers the standard and ubiquitous price-of-risk vs price-volatility relationship:
@@
Remark 10.
The proportionality factor characterises ex-ante pricing through an intuitive and direct metric of asset-pricing (Remark 7). How its level comes about is a matter of economic theory (e.g. risk-aversion). Regardless, our derivation links it firmly to ex-ante beliefs, making the above relevant to bias studies and measurement under model-uncertainty.
@@ 3.1 The Joint-Hypothesis Issue
To focus on main issues, as in Section 2.2.2, consider usual-case assets under model-risk $B$ with $B$-independent $B$-sure risk-premia $\{\check{rp}(t)\}$. Let $\{\mathbf{p}^{B}_{t}\}$ be the $\{\mathbf{F}_{t}\}$-adapted true conditional probabilities of model-risk outcomes, given the true and known $B$-sure laws $\mathbf{W}^{T}_{B}$ and the true but potentially unknown unconditional law $\mathbf{p}^{B}_{0}$ of model-risk $B$. Its ex-post price-drift to be realised on average $\{rp_{t}\}:=\{\Sigma_{B}\mathbf{p}^{B}_{t}{Y^{B}_{t}}\}-\{S_{t}\}$ satisfies:
@@ 3.2 Bias and Risk-Pricing Parameters under Model-Uncertainty
The triggering events may or may not follow known laws. The process of how/when they arise is ignored, assuming that the average time between model crises, say some $\lambda^{-1}\gg 1$, is so long that each can be considered one-off. The period of interest is right after the arrival of a trigger but long before the next, so $t\ll\lambda^{-1}$, and long before horizon/maturity, so $t\ll T$.
@@
If the change-risk is diversifiable, so unpriced, then $K=1$; if it is classical, so its law $\mathbf{p}^{1}_{0}$ is known, then $\rho=1$. Parameter $\rho$ stems from model-uncertainty and is thus unknown ex-ante.
@@
3.3 An Ideal Model-Uncertainty Driven Market
Consider a model-uncertainty driven market, of a large number of assets, each facing some change-risk $B$, whose nature may vary from asset to asset (e.g. new CEO, tort, war) but whose true $B$-sure laws are known. Let the $B$-risks be priced (i.e. correlated with the market’s stochastic-discount-factor), for their effects on aggregate value/volatility/cross-sectional dispersion [^6] Dispersion within members of an equity index, as a priced factor, has been linked to market uncertainty, the business cycle and structural shifts (e.g. Demirer and Jategaonkar (2013), Kolari et al. (2021)).. Further, let them be economically priced, so that by Proposition 1 their pricing is canonical with respect to some reference inferential belief process of the relevant $B$-risk.
Finally, let the realisation and detection of individual change-outcomes be stochastic and independent (even for macroeconomic triggers such as war or recession). Note that such a market, when seen as a whole, is an asset with non-binary model-uncertainty; this is so even when there is only one binary macroeconomic model-risk trigger.
For brevity, let the unconditional all-cause probability of positive/negative economic change be symmetrical: at any time half of all potential changes, of any origin, are positive/negative (unconditional asymmetry is addressed in Remark 14&15). Let ’pricing intention’ $K$ ((39)) and ’apparent bias’ $\rho$ ((38)) be indifferent to the economic direction and origin of change-risk, and the size of $K$, $\rho$ and risk-impact $S^{\Delta}$, uniform across the assets. Such uniformity in practice amounts to asset dispersions being driven by differences in the paths of model-risk inference rather than in any other aspect (Remark 9). Similarly, let all risk-inference be based on log-LR processes that are homogenous in time and uniform across.
@@ 3.4.1 Momentum Effect and Status Quo Bias
Momentum applies also to $\rho\ll 1$: past excess due to RNE beliefs $\{\Pi^{1}_{t}\}$ falling from $\Pi^{1}_{0}\approx 1$ to $v<\Pi^{1}_{0}$ means an average future excess of the same sign; indeed, label-switching defines a new parameter $\rho^{-1}>1$. Momentum thus vanishes only when $\rho=1$ (e.g. when there is no model-uncertainty).
@@ 3.4.2 Low-Risk Effect and Status Quo Bias
At $\rho=1$ the above is exact, by which $v^{\sigma}_{max}=\frac{1}{2}$, where price-volatility peaks, thus confirming ’high risk high reward’ in the absence of model-uncertainty or bias. There is a low-risk effect otherwise: peak-reward occurs at $v^{\sigma}_{max}\approx\rho^{-1}$, while peak volatility ($v=\frac{1}{2}$) is poorly rewarded ex-post: $rp^{\sigma}(\frac{1}{2})=\frac{S^{\Delta}}{2}\mathcal{O}(\frac{K-1}{\rho}-1)$.
@@ 4 Summary and Discussion
We study the risk-neutral equivalent formulation of model-risk pricing. Explicit price dynamics under binary model-risks with potentially unknown laws are derived.
The results show that model-risk can be a main source of excess value, with a key role in Status Quo Bias and anomalies such as Momentum and Low-Risk. They offer an integrated and practicable approach to these highly relevant and interconnected topics.
We arrive at the above by making use of the informational constraint imposed by ’usual’ economic asset-pricing (Remark 1 and Definition 1), to show that viable asset-pricing under model-risk may take just one canonical form (Lemma 1 and Proposition 1), which leads to a unique and intuitive model-risk pricing formula with convenient properties (Corollary 1).
The sole risk-pricing parameter is associated with a conserved quantity of inferential dynamics and is a product of ’intention’ (preference) and ’mistake’ (difference between objective and subjective laws). Ex-ante (’intended’) pricing, classical or with ambiguity-aversion/robust-control, can be captured by our parameter $K\in[1,\infty)$, (39), with $K=1$ corresponding to ’no model-risk discounts’. Its derivation reveals that it sets the intended gain-to-loss ratio of the asset at peak model-risk, suggesting $K\in[1,2)$ in competitive markets (Remark 7).
@@
The above is interesting also in light of the inertia axiom of Bewley (2002), which says, in effect, that to update an existing decision by incorporating a new state subject to Knightian uncertainty, the status quo is kept unless an alternative that is better in all scenarios exists. For our model-risk, it suggests inertial pricing, one that is piece-wise free of model-risk pricing, jumping between alternative models when appropriate, despite data that can resolve model-risk via regular inference. Asset-prices under max-min utility already have this characteristic, albeit with a built-in pessimism absent in inertial pricing. Both are viable in our setup, as ’models’ are represented by probability laws that are equivalent before any risk-resolution (that is predictable). The theoretical and practical aspects of this topic will be examined in a follow-up study.

### v8

@@ 1.3 The Economic Decomposition of Risk-Pricing under Model-Risk
where $\check{\mathbf{E}}_{{B}}[\cdot|\cdot]$ denotes expectations under its $B$-sure RNE measure $\check{\mathbf{W}}^{T}_{B}\sim{\mathbf{W}}^{T}_{B}$, leading to RNE outlook $\check{y}_{B}(T-t,t)={\mathbf{E}}[\int_{t}^{T}\check{r}_{B,s}ds]$ as in (2), and $B$-sure risk-premium $RP^{B}(t)$ satisfying: $\forall t<T$,
@@
where (14) ensures consistency with asset-property Item-3, Section 1.1, while (15), the internal consistency of the pricing scheme under model-risk $B$.
@@
$k^{A}_{t}:=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}\geq 0,$ (19)
@@
In ambiguity-aversion settings, the primary utility function, enforcing risk-aversion, and the secondary-utility/variational-control, enforcing uncertainty-aversion, can be dialled independently and arbitrarily close to neutrality (linearity), to achieve the above separation. In classical and some parameter-learning settings, with a single utility, the above obtains by replacing payoffs with their model-sure expectations perturbed by suitably small noise.
@@
Asset-pricing under model-risk $B$ is said to be usual if its price-of-$B$-risk $k^{A}_{t}$, (19), is a time-homogenous twice-differentiable function of its model-risk belief [^3] Economic risk-pricing, as the solution to some optimisation scheme, is measurable to the natural filtration of its input $(\{Z_{t}\},\{\pi^{B}_{t}(\mathbf{D}_{t})\})$; so it is at any time $t$ a function of $(Z_{t},\pi^{B}_{t}(\mathbf{D}_{t}))$ (Proposition 4.9 of Breiman (1992)). Any explicit dependence on $Z_{t}$ or time, if at all, is presumed weak or non-existent, at least for $t\ll T$. $\pi^{B}_{t}$, $\forall t\ll T$.
@@ 2.2.1 The 1-Parameter Price-of-Model-Risk Formula
The difference process $\{{\pi}^{+}_{t}\}-\{{\Pi}^{+}_{t}\}=\{k^{\Pi}_{t}\sigma^{\pi}_{t}\}$ has a fixed sign. Unless the model-risk is consistent (see Item-3, Section 1.1, and (14)), viable model-risk premia $\{B\_RP_{t}\}$ (see (18)) must be negative sometimes and so correspond to no known economic risk-pricing theories.
@@ 2.2.2 Connection with Familiar Price-Dynamics
For better focus, let any $B$-sure risk-premia be indifferent to $B$-value henceforth, to be written as $\check{rp}(t)={\mathbf{E}}[\int_{t}^{T}\check{R}_{s}ds]>0$, with $\check{R}_{t}$ denoting the $B$-sure risk-premium drift at any $t<T$; note then $Y^{\Delta}(t)=S^{\Delta}(t)$ as a result, Canonical pricing (22) becomes:
@@
$dS_{t}(B)=dY^{-}_{t}(B)+S^{\Delta}(t)d\Pi^{+}_{t}(B)-r^{Z\Delta}_{t}\Pi^{+}_{t}(B)dt+\check{R}_{t}dt.$ (25)
By setup (Section 1.1) and inference dynamics (A.2-A.3), process $\{d\Pi^{+}_{t}\}$ has two sources of noise, one from $\{Z_{t}\}$ ((1)), given by $\{\sigma^{lZ}_{t}dw^{Z}_{t}\}$, with $\sigma^{lZ}_{t}:=r^{Z\Delta}_{t}/\sigma^{Z}_{t}\ll 1$ (’signal-to-noise’), one from independent $B$-informative data $\{D_{t}\}$ ((3)), modelled likewise, which can be absorbed through $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$ and $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$. Therefore:
@@
$\mu^{\Pi,\pi}_{t}:=\Sigma_{B}{\pi}^{B}_{t}{\mu^{\Pi,B}_{t}}=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot(\pi^{+}_{t}-\Pi^{+}_{t})=(\sigma^{\Pi}_{t}\sigma^{l}_{t})^{2}\cdot k^{\Pi}_{t}\sigma^{\pi}_{t}\geq 0;$ (29)
@@
$+$ $\big{[}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{t}(B)\big{)}\cdot r^{Z\Delta}_{t}dt\big{]}$ (31)
@@
Inference dynamics (32) usually dominate $B$-sure dynamics (30) and model-drift (31), as $\check{R}_{t}\ll S^{\Delta}(t)$ and $r^{Z\Delta}_{t}\ll S^{\Delta}(t)$. Indeed, during gaps between payoff updates, which exist in reality, our asset-price equation of motion reduces to (32), essentially the sequential test process of Peskir and Shirayev (2006). Further, it often holds that $(\sigma^{l}_{t})^{2}\approx(\sigma^{lD}_{t})^{2}$ (relative variance $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}\ll 1$), so that model-risk inference, not tangible value, is the main driver of price moves and dispersions among the assets affected. Finally, under model-risk $B$, even if unpriced (so the inference-induced drift (29) is nil), asset-price volatility rises regardless, unless asset-pricing is already $B$-sure (i.e. $(\sigma^{\Pi}_{t})^{2}=0$).
@@
Expansion in (K1)(K-1), given Remark 8, recovers the standard and ubiquitous relationship:
@@
The proportionality factor characterises risk-pricing via an intuitive metric, the intended gain-to-loss (Remark 8). How its level comes about is up to economic theories (e.g. risk-aversion). Our derivation, linking it to ex-ante beliefs, makes it relevant to bias studies.
@@ 3.1 The Joint-Hypothesis Issue
Take as in Section 2.2.2 usual-case assets with model-risk $B$ and $B$-indifferent $B$-sure risk-premia $\{\check{rp}(t)\}$, for better focus. Let $\{\mathbf{p}^{B}_{t}\}$ be the $\{\mathbf{F}_{t}\}$-adapted true conditional probabilities of $B$-values, given the true, known, $B$-sure laws $\mathbf{W}^{T}_{B}$, and the true, potentially unknown, unconditional law $\mathbf{p}^{B}_{0}$ of $B$. Its ex-post average price-drift $\{rp_{t}\}:=\{\Sigma_{B}\mathbf{p}^{B}_{t}{Y^{B}_{t}}\}-\{S_{t}\}$ satisfies:
@@ 3.2 Bias and Risk-Pricing Parameters under Model-Uncertainty
We ignore the law governing triggering events by assuming that the average time between them, say $\lambda^{-1}\gg 1$, is so long that each is considered one-off. The period of interest is just after an arrival, long before the next, and long before horizon, so $t\ll\lambda^{-1}$ and $t\ll T$.
@@
If the change-risk is diversifiable, so unpriced, then $K=1$; if it is classical, so its law $\mathbf{p}^{1}_{0}$ is known, then $\rho=1$. Parameter $\rho$ is born of model-uncertainty and thus unknown ex-ante.
@@
3.3 An Ideal Market Driven by Model-Uncertainty
Consider an ideal market, of a large number $N\gg 1$ of usual-case assets, each facing a change-risk $B$, whose nature may vary from asset to asset (e.g. new CEO, tort) but whose $B$-sure laws are known. Let them be priced-risks (i.e. correlated with the stochastic-discount-factor) of the market, as they can elevate total volatility and cross-sectional dispersion [^6] Dispersion among members of an equity index, as a priced factor, has been linked to market uncertainty, the business cycle and structural shifts (e.g. Demirer and Jategaonkar (2013), Kolari et al. (2021))., irrespective of their impact on total value (see (32) and Remark 10). Let them be economically priced, so that Section 2 applies (to individual asset processes).
Note that the market as a whole has a non-binary model-risk: even if there is just one binary trigger (e.g. war, recession), there can be as many as $2^{N}$ market models. Let the realisation and detection of each individual change-outcome be stochastic and independent, even for a macroeconomic trigger.
For brevity, let the unconditional all-cause probability of positive/negative economic change be symmetrical: at any time half of all potential changes, of any origin, are positive/negative (unconditional asymmetry is addressed in Remark 14&15). Let ’pricing intention’ $K$ ((39)) and ’apparent bias’ $\rho$ ((38)) be indifferent to the economic direction and origin of change-risk, and the size of $K$, $\rho$ and risk-impact $S^{\Delta}$, uniform across the assets. Such uniformity in practice amounts to asset dispersion being driven primarily by differences in the evolutions of model-risk inference rather than in any other aspect (Remark 10). Similarly, let all risk-inference be based on log-LR processes that are homogenous in time and uniform across.
@@ 3.4.1 Momentum Effect and Status Quo Bias
Momentum applies also to $\rho\ll 1$: past excess due to RNE beliefs $\{\Pi^{1}_{t}\}$ falling from $\Pi^{1}_{0}\approx 1$ to $v<\Pi^{1}_{0}$ means an average future excess of the same sign; indeed, label-switching defines a new parameter $\rho^{-1}>1$. Momentum vanishes only if $\rho=1$ (no bias or model-uncertainty).
@@ 3.4.2 Low-Risk Effect and Status Quo Bias
At $\rho=1$ it is exact and $v^{\sigma}_{max}=\frac{1}{2}$ (peak volatility), thus confirming ’high risk high reward’ in the absence of model-uncertainty or bias. Otherwise a low-risk effect applies: peak-reward occurs at $v^{\sigma}_{max}\approx\rho^{-1}$, and peak volatility ($v=\frac{1}{2}$) is poorly rewarded ex-post: $rp^{\sigma}(\frac{1}{2})=\frac{S^{\Delta}}{2}\mathcal{O}(\frac{K-1}{\rho}-1)$.
@@ 4 Summary and Discussion
We propose a risk-neutral equivalent formulation of model-risk pricing. Explicit price dynamics under binary model-risks with potentially unknown laws are derived. One implication is that model-risk can be a main source of excess value, with a key role in Status Quo Bias and anomalies such as Momentum and Low-Risk. Our approach offers an integrated and practicable framework to study these highly relevant and interconnected subjects.
We arrive at the above by making use of the informational constraint imposed by ’usual’ economic asset-pricing (Remark 1 and Definition 1), to show that viable asset-pricing under model-risk may take just one canonical form (Lemma 1 and Proposition 1), which leads to a unique and intuitive model-risk pricing formula with convenient properties (Corollary 1). The sole risk-pricing parameter derives from a conserved quantity of inference dynamics and is a product of ’intention’ (preference) and ’mistake’ (difference of objective and subjective laws).
Ex-ante (’intended’) pricing, classical or with ambiguity-aversion/robust-control, can be captured by our parameter $K\in[1,\infty)$, (39), with $K=1$ corresponding to ’no model-risk discounts’. Its derivation reveals that it sets the intended gain-to-loss ratio of the asset at peak model-risk, suggesting $K\in[1,2)$ in competitive markets (Remark 8).
@@
The above is interesting also in light of the inertia axiom of Bewley (2002), which says, in effect, that to update an existing decision by incorporating a new state subject to Knightian uncertainty, the status quo is kept unless an alternative that is better in all scenarios exists. For our model-risk, it suggests inertial pricing, one that is piece-wise free of model-risk pricing, jumping between alternative models when appropriate, despite data that can resolve model-risk via regular inference. Asset-prices under max-min utility already have this characteristic, albeit with a built-in pessimism absent in inertial pricing. Both are viable in our setup, as ’models’ are represented by probability laws that are equivalent before any risk-resolution. The theoretical and practical aspects of inertial pricing will be examined in a follow-up study.